# Quality Time: Carbon-Aware Quality Adaptation for Energy-Intensive Services

### v1

@@ 4.3. Dynamic QoR Adaptation for Best-Effort Users
Note, that we solve the optimization over $i=0,\dots,I-1$ (not just $i\geq\alpha$) as we must aggregate the emissions from all intervals in Equation (14). This has no performance implication, as we still only optimize the variables $(D^{i},A^{i})_{i=\alpha}^{I-1}$. Past values in $D^{i}$ and $A^{i}$ are fixed.
@@ 5. Experiments
We assess the potential of carbon-aware QoR adaptation by simulating the execution of an LLM inference service over the entire year 2023 using two real-life request traces and in ten different regions.
All experiments were conducted under the same conditions on an HPC cluster, with 16 cores allocated to each job. We use Gurobi (Gurobi Optimization, LLC, 2024) for solving the optimization problems.
@@ 5.1. Experimental Setup
As average carbon intensity is usually reported hourly, we consider $\Delta=1$ hour.
Scenario
@@
We model machine performance using recent benchmarks of vLLM on a EC2 p4d.24xlarge for inference, which state a throughput of 11.57 requests per second for LLaMA 3.1 8B and 5.05 requests per second for LLaMA 3.1 70B (vLLM Team, [n. d.]).
Request traces
We simulate realistic request patterns using two hourly request traces from the Wikipedia Pageview statistics (Wikimedia Foundation, [n. d.]), as shown in Figure 5. In particular, we use
  • Wiki (en): the English Wikipedia with $63.4e6\pm 0.8e$ 6 requests per hour and spikes of up to $616.4e$ 6 requests per hour. In our scenario, this results in 80-200 active machines throughout the year.
  • Wiki (de): the German Wikipedia with $54.7e5\pm 2.8e$ 5 requests per hour, with spikes of up to $517.6e$ 5 requests per hour. This results in only 10-25 active machines, emphasizing the importance of discrete scheduling decisions.
Figure 5. The request traces differ notably, as the German Wikipedia is primarily accessed from a single timezone, while the English Wikipedia sees traffic from around the globe.
Carbon intensity regions
We analyze the potential in ten different regions:
@@
  • In Australia, we consider Queensland (AUS-QLD)
All carbon intensity data was provided by ElectricityMaps (Maps, 2024). We normalize all time zones to align patterns with the time of day in the request traces. Temporal variability of hourly carbon intensities are shown in Figure 1.
5.2. Impact of QoR on Energy Demand
We first analyze the absolute yearly emissions generated by our different scenarios without carbon-aware QoR adaption. For this, we solve Equation (9) for both request traces, across all 10 regions, $\text{QoR}_{\text{target}}=\{0,0.1,...,1\}$, and $\gamma=1$ hour. The results reflect emission savings driven purely by energy demand, i.e., by serving more or less requests on Llama 3.1 70B. We optimize until the objective is within 0.1 % of its optimal solution or after 10 minutes. All results are within 0.8 % of the optimal solution.
Figure 6 shows the overall emissions for Wiki (en). As expected, we observe very large differences across different regions. For example, Sweden’s energy mix relies almost entirely on renewable energy and nuclear power, making it one of the least carbon-intensive regions globally. Serving all requests with the 80B model results in just 245 tCO2, almost 27 times less than in Poland (4681 tCO2), where coal is the dominant energy source. As expected, the resulting emissions scale linearly with $\text{QoR}_{\text{target}}$. Results are equivalent for Wiki (de), but smaller by a factor of around 7.5.
Figure 6. Annual operational carbon emissions for Wiki (en) at different $\text{QoR}_{\text{target}}$ for all 10 regions.
Figure 7. Additional carbon savings resulting from increased validity periods, relative (top) and absolute (bottom) for $\text{QoR}_{\text{target}}=0.5$. Bigger validity periods yield larger gains due to increased flexibility.
5.3. Upper-Bound Potential of Carbon-Aware QoR Adaptation
@@
Carbon Savings Beyond Energy Efficiency
As define scenarios with validity periods of $\gamma=1$ hour as baselines. For these scenarios, emissions are directly correlated with energy demand as there is no flexibility for carbon-aware QoR adaptation. Figure 7 reports the relative and absolute additional savings for $\text{QoR}_{\text{target}}=0.5$ and different validity periods with respect to these baselines.
@@
Increasing $\gamma$ to 8 hours yields only limited additional savings—less than 3 % across all settings. This is because average carbon intensity shows limited short-term variation, leaving little room for optimization. However, starting at $\gamma=24$ hours, we can see more notable potential of 5-8 % in some regions. The characteristics of carbon intensity data significantly influence the expected gains. For instance, while California’s carbon intensity primarily fluctuates on a daily basis due to its heavy reliance on solar energy, Germany experiences more complex patterns with daily, weekly, and seasonal variations (Wiesner et al., 2021). As a result, extending $\gamma$ beyond 24 hours has minimal impact in California, whereas in Germany, larger $\gamma$ values lead to considerable improvements. In contrast, regions like Sweden, New York, and PJM exhibit very little temporal variation in carbon intensity, which limits the potential for carbon-aware approaches.
Impact of $\text{QoR}_{\text{target}}$
Figure 8. Additional relative savings for different $\text{QoR}_{\text{target}}$ in Germany and California.
@@
Figure 8 shows the potential for relative savings across different $\text{QoR}_{\text{target}}$. When $\text{QoR}_{\text{target}}$ is 0, the optimizer will serve all requests using LLaMA 3.1 8B, leaving no flexibility for additional savings. Likewise, when $\text{QoR}_{\text{target}}$ is 1, all requests must be served by LLaMA 3.1 70B. Values around $\text{QoR}_{\text{target}}=0.5$ offer the greatest flexibility for carbon-aware QoR adaptation, resulting in the highest potential for carbon savings.
Figure 9. Additional carbon savings under realistic conditions. Our online approach maintains additional savings within one percentage point across most experiments.
Impact of validity periods
(a) $\beta=\{1,4,8,24\}$ hours for smaller validity periods.
(b) $\beta=\{1,3,7\}$ days for larger validity periods.
@@
Figure 10. Considering all time intervals of length $\beta$ throughout the year, what proportion of these intervals have a QoR below the threshold? We analyze the CDF for Wiki (en) in Germany for $\text{QoR}_{\text{target}}=0.5$.
While increasing $\gamma$ enhances potential for carbon savings, it also raises the risk of prolonged periods with low QoR. Service providers can adjust the range of QoR using SLO and SLG to ensure that a certain fraction of requests can always be served from a specific quality tier. However, limiting the fraction of requests available for QoR adaptation lowers its potential impact.
Figure 10 illustrates the probability of prolonged periods with low QoR under different validity period lengths. To to provide an example of how to interpret the results: Figure 10(a) shows that when optimizing for a $\text{QoR}_{\text{target}}=0.5$ with validity periods of $\gamma=8$ hours, no 8-hour interval ($\beta$) has a QoR below 0.5, while all intervals maintain a QoR of at least 0.5. However, around 10 % of all 4-hour intervals fall below a QoR of 0.25.
We observe, that for validity periods of $\gamma=3$ months, users can experience several consecutive days of low QoR. Consequently, we conclude that very large validity periods of more than a month are not practical in real-world scenarios.
5.4. Potential under Realistic Conditions
@@
We evaluate the practicability of carbon-aware QoR adaptation using our proposed online optimization under real forecasts and near-optimal MILP solutions.
Setup
We simulate request forecasts using Prophet (Taylor and Letham, 2017), a state-of-the-art forecasting model that fits daily, weekly, and yearly seasonalities. In particular, we fit a model every day at midnight using 3 years of historical data to forecast the rest of the year. The predictions exhibit realistic errors: For Wiki (en), 24-hour forecasts have an average MAPE of $13.9\pm 8.4$, while for Wiki (de), the error is $32.1\pm 15.3$.
For long-term carbon intensity forecasts, we predict the entire year of 2023 as of January 1st using Prophet, fitted on three years of historical data. We simulate short-term forecasts of up to 4 days by adding Gaussian noise to the actually observed carbon intensity data to match the MAPEs reported by the state-of-the-art forecasting model CarbonCast (Maji et al., 2023) for each region (see Appendix A). Short-term forecasts are updated daily at midnight. We interrupt long-term optimizations after 30s and short-term optimizations after 10s, to demonstrate performance under near-optimal solutions.
Furthermore, we limit this analysis to daily, weekly, and monthly validity periods, as Section 5.3 showed that validity periods shorter than a day offer minimal potential for carbon savings, while periods longer than a month can lead to extended durations of low QoR.
Results
Figure 9 shows the performance gap of our online optimization compared to the theoretical upper bound established in Section 5.3. Across all load traces and regions, the resulting emissions savings are generally within one percentage point of the optimal solution. In regions with already limited upper-bound potential, the expected savings become more marginal. However, in most regions, our approach consistently maintains additional savings of 5–10 %.
Within the stated time limits, 95 % of all long-term optimizations achieve an optimization gap $<2.1\,\%$, while 95 % of all short-term optimizations reach a gap $<0.3\,\%$.
5.5. Outlook of Yearly Emissions
(a) Yearly emission outlook at three discrete points in time.
(b) Yearly emission outlook, performed weekly for $\text{QoR}_{\text{target}}=0.5$.
Figure 11. Projected yearly emissions (blue) based on long-term forecasts for different $\text{QoR}_{\text{target}}$. The interval shows the range between best- and worst-case forecasts of $\widehat{R}$, while red indicates emissions already incurred.
Long-term optimizations become computationally cheap when accepting near-optimal solutions. This allows us to generate a broad spectrum of solutions for varying $\text{QoR}_{\text{target}}$ and request forecasts $\widehat{R}$, which may represent, for instance, worst-case or best-case scenarios.
Setup
@@ 5.5. Outlook of Yearly Emissions
We demonstrate emission outlooks by executing the online approach on the Wiki (en) traces in California. For simulating best-case and worst-case request forecasts, we use the lower and upper bound of the 80 % prediction intervals provided by Prophet, respectively. Each week, we solve 33 long-term optimizations, covering the three forecasts and $\text{QoR}_{\text{target}}=\{0,0.1,...,1\}$.
Results
@@
Figure 11(b) shows that the initial estimate was overstated, although still in the range of the lower bound It depicts the weekly estimates for annual emissions when aiming for $\text{QoR}_{\text{target}}=0.5$. We observe that the upper and lower bound, defined by the worst-case and best-case forecasts, narrow over time as actual emissions accumulate.
5.6. Dynamic QoR Adaptation for Best-Effort Users
Figure 12. Daily QoR for California, Poland, and Sweden. Our adaptive approach provides the most consistent QoR throughout the year.
Setup
We illustrate the dynamic adaptation of $\text{QoR}_{\text{target}}$ introduced in Section 4.3 on Wiki (en) in California, Poland, and Sweden to cover a spectrum of differently behaving carbon intensity traces: California exhibits high variability due to many solar deployments, Poland has relatively little variation due to reliance on fossil fuels, and Sweden usually has very low carbon intensity which occasionally spikes when demand needs to be met through cross-border imports or natural gas. For each region, we define the annual budget $B^{\text{annual}}$ so that the ideal $\text{QoR}_{\text{target}}=0.50$.
We compare our approach to two greedy baselines:
@@ 5.6. Dynamic QoR Adaptation for Best-Effort Users
    Greedy (constant) divides the annual budget into equally sized hourly budgets for which it solves Equation (13). Any remaining or overspent budget is evenly distributed across the remaining time steps.
@@
    Greedy (weighted) solves Equation (13), too, but weights the hourly budgets based on forecasted load and carbon intensity. Every 24 hours, as forecasts are updated, weightings get adjusted.
Results
Figure 12 illustrates the daily QoR of our online approach compared to both greedy baselines over time. Since Greedy (constant) does not account for fluctuations in requests or carbon intensity, the provided QoR fluctuates significantly throughout the year. Greedy (weighted) delivers more consistent QoR for both Poland and Sweden, though it starts very conservatively in both regions, leading to near-perfect QoR at the year’s end due to unused budget. In California, where forecasts are less accurate, this baseline overspends early in the year, resulting in days with $\text{QoR}=0$ for most of the year’s end. Our approach also starts conservatively, as the forecast initially overestimates annual emissions (see Figure 11). However, it delivers the most consistent service across all experiments.

### v2

@@ 4.2. Automatic QoR Adaptation
Note, that we solve the optimization over $i=0,\dots,I-1$ (not just $i\geq\alpha$) as we must aggregate emissions over all intervals in (16). This has no performance implication, as we still only optimize the variables $(D^{i},A^{i})_{i=\alpha}^{I-1}$. Past values in $D^{i}$ and $A^{i}$ are fixed.
@@ 5. Experiments
We assess the potential of carbon-aware QoR adaptation by simulating the execution of an LLM inference service over the year 2023 using 8 real and synthetic request traces in 10 different regions.
We use Gurobi (Gurobi Optimization, LLC, 2024) to solve optimization problems. All simulations were conducted under the same conditions on an HPC cluster, with 16 cores allocated to each job.
Table 2. Upper-bound potential and carbon savings under realistic conditions in % for all request traces and regions at $\gamma=1$ week and $\text{QoR}_{\text{target}}=0.5$. Note, that other $\text{QoR}_{\text{target}}$ can lead to even better results (Figure X).
Region Artificial Real Synthetic Mean
Static Random Wiki (en) Wiki (de) Taxi Cell B Cell D Cell F
NL 8.2 / 7.4 8.5 / 6.3 7.6 / 6.5 10.2 / 9.3 11.3 / 8.7 7.8 / 6.7 7.6 / 6.6 8.0 / 7.0 8.7±1.3 / 7.3±1.0
CISO 7.0 / 6.8 7.5 / 6.0 7.5 / 7.0 10.0 / 9.4 11.3 / 9.2 6.6 / 6.1 6.3 / 5.9 6.8 / 6.4 7.9±1.7 / 7.1±1.3
ES 7.1 / 6.3 7.3 / 5.2 6.5 / 5.6 9.4 / 8.4 10.4 / 8.0 6.6 / 5.6 6.6 / 5.7 6.9 / 5.9 7.6±1.4 / 6.3±1.1
AU-QLD 7.9 / 7.8 8.4 / 7.0 7.1 / 6.5 6.1 / 5.8 8.8 / 6.9 7.6 / 7.0 6.9 / 6.4 8.0 / 7.5 7.6±0.8 / 6.9±0.6
DE 6.3 / 5.4 6.7 / 4.5 5.7 / 4.6 8.7 / 7.6 9.7 / 6.9 5.9 / 4.8 5.8 / 4.8 6.1 / 5.1 6.9±1.4 / 5.5±1.1
PL 4.1 / 3.8 4.5 / 2.9 3.5 / 2.9 6.7 / 6.2 7.5 / 5.7 3.8 / 3.3 3.7 / 3.2 4.1 / 3.5 4.7±1.4 / 3.9±1.2
ERCOT 4.2 / 3.1 4.6 / 2.6 3.6 / 2.4 6.3 / 5.1 7.4 / 5.8 3.9 / 2.8 3.7 / 2.5 4.1 / 3.0 4.7±1.3 / 3.4±1.2
SE 2.7 / 2.4 3.0 / 1.6 2.0 / 1.5 4.9 / 4.5 6.1 / 5.4 2.3 / 1.9 2.2 / 1.7 2.6 / 2.2 3.2±1.4 / 2.6±1.4
NYISO 2.3 / 1.8 2.7 / 1.1 1.7 / 1.1 4.6 / 4.3 5.6 / 4.6 1.9 / 1.5 1.8 / 1.3 2.2 / 1.7 2.8±1.4 / 2.2±1.3
PJM 1.8 / 1.7 2.3 / 0.8 1.3 / 1.0 4.1 / 4.0 5.5 / 4.3 1.6 / 1.4 1.4 / 1.1 1.8 / 1.6 2.5±1.4 / 2.0±1.3
@@ 5.1. Experimental Setup
As ACI is usually reported hourly, we define $\Delta=1$ hour.
@@
Request traces.  For our analysis, we require 4 years of hourly request traces: 2020-2022 for fitting forecasting models and 2023 for the analysis itself. As such data are not publicly available for LLM inference servies, we show the effectiveness of our method across a diverse set of other real and synthetic request traces:
  • Artificial: A Static trace that assumes a constant stream of hourly requests, and a Random trace, where we sample hourly requests from a normal distribution.
  • Real: Two real-life traces representing the English and German Wikipedia Pageview statistics (Wikimedia Foundation, [n. d.]), Wiki (en) and Wiki (de), and a trace based on the hourly taxi trips in New York City (Taxi and Commission, 2025), called Taxi.
  • Synthetic: Three generated traces, called Cell B, D, and F, which are based on instance event traces of Borg cluster cells at Google (Google, 2019).
All request trace datasets and our forecast methodology are explained in detail in Appendix B.
Carbon intensity traces.  We evaluate the potential for ten globally distributed regions:
@@
  • In Australia, we consider Queensland (AU-QLD)
All carbon intensity data was provided by ElectricityMaps (Maps, 2024). We normalized all time zones to align the time of day with the request traces. The temporal variability of regions is depicted in Figure 1. Our forecast methodology is explained in Appendix C.
5.2. Absolute Emissions
We first analyze the absolute annual emissions without carbon-aware QoR adaption, i.e. for $\gamma=1$ hour. The results reflect emission savings driven purely by increased energy demand from serving more requests on Llama 3.1 70B ($\text{QoR}_{\text{target}}\to 1$). Figure 4 shows the annual emissions for Wiki (en)—the trend is equivalent for all datasets. We observe very large differences across different regions: For example, Sweden’s energy mix relies almost entirely on renewable energy and nuclear power, making it one of the least carbon-intensive regions globally. Serving all requests with the 70B model results in just 245 tCO2, almost 27 times less than in Poland (4681 tCO2), where coal is the dominant energy source. As expected, the resulting emissions scale linearly with $\text{QoR}_{\text{target}}$.
Figure 4. Annual operational carbon emissions for Wiki (en) at different $\text{QoR}_{\text{target}}$ without carbon-aware QoR adaption.
Figure 5. Additional carbon savings resulting from increased validity periods, relative (top) and absolute (bottom) for $\text{QoR}_{\text{target}}=0.5$. Bigger validity periods yield larger gains.
5.3. Upper-Bound Potential
@@
Carbon savings beyond energy efficiency.  In blue (left), Table 2 presents the relative additional carbon savings for $\text{QoR}_{\text{target}}=0.5$ at $\gamma=1$ week compared to $\gamma=1$ hour, where emissions directly depend on energy demand. We observe that through carbon-aware QoR adaptation, emissions drop by around 8 % in many regions. While request trace patterns exhibit little difference in potential, traces with a low volume of requests (Wiki (de) and Taxi) show increased potential, due to the higher impact of discrete scheduling decisions. For example, the volume of requests in Wiki (de) results in 10–25 active machines at a time, while Wiki (en) requires 80–200 active machines. In experiments with even fewer machines we managed to gain relative savings of more than 40 %—however, we do not analyze such cases further, as we focus on high-impact scenarios with many machines.
@@
Impact of validity periods.  Figure 5 reports the relative and absolute savings for $\text{QoR}_{\text{target}}=0.5$ and different validity periods. Increasing $\gamma$ to 8 hours yields only limited additional savings—less than 3 % across all settings. This is because average carbon intensity shows limited short-term variation, leaving little room for optimization. However, starting at $\gamma=24$ hours, we see a more notable potential of 5-8 % in some regions. Regional characteristics significantly influence the expected gains: For instance, while California’s carbon intensity primarily fluctuates on a daily basis due to its heavy reliance on solar energy, Germany exhibits more complex patterns with daily, weekly, and seasonal variations (Wiesner et al., 2021). As a result, extending $\gamma$ beyond 24 hours has minimal effect in California, while in Germany, larger $\gamma$ values lead to considerable improvements. In contrast, regions like Sweden, New York, and PJM exhibit very little temporal variation in carbon intensity.
Impact of premium users.  Figure 6 illustrates the loss of potential with a growing fraction of premium users at $\gamma=1$ week and $\text{QoR}_{\text{target}}=0.5$. We observe a slightly sub-linear decrease with a certain offset which depends on the overall request volume and, hence, the impact of discrete scheduling decisions. For reference, ChatGPT reached over 200 million active users in 2024 (Reuters, 2024), with fewer than 10 million paying customers (Tarantola, 2024). The resulting 5 % premium users decrease the potential savings by 5.6 % on average.
Figure 6. Impact of premium users on the potential savings.
@@
Impact of $\text{QoR}_{\text{target}}$.  Figure 7 shows the potential for relative savings across different $\text{QoR}_{\text{target}}$. At $\text{QoR}_{\text{target}}=0$, we serve all requests using LLaMA 3.1 8B, leaving no flexibility for additional savings. Likewise, when $\text{QoR}_{\text{target}}$ is 1, all requests must be served by LLaMA 3.1 70B. A $\text{QoR}_{\text{target}}$ around 0.5 offers the greatest flexibility for carbon-aware QoR adaptation, resulting in the highest potential for carbon savings.
Figure 7. Additional relative savings for different $\text{QoR}_{\text{target}}$.
@@
Figure 8. CDF for Wiki (en) in Germany at $\text{QoR}_{\text{target}}=0.5$: What proportion of time intervals of length $\beta=\{1,3,7\}$ days have a QoR below the threshold?
Periods of low service quality.  While increasing $\gamma$ enhances potential for carbon savings, it also increases the probability of prolonged periods with low QoR, illustrated in Figure 8. For example, when optimizing for $\text{QoR}_{\text{target}}=0.5$ at $\gamma=1$ week, no 1-week interval has a QoR below 0.5 (right). However, 10 % of all daily intervals exhibit a $\text{QoR}_{\text{target}}<0.48$ (left). We observe that for $\gamma=3$ months, users can experience multiple consecutive days of low QoR. Consequently, very large validity periods of more than a month are not practical in real-world scenarios.
5.4. Evaluation under Realistic Conditions
@@
We evaluated the practicability of carbon-aware QoR adaptation using the proposed multi-horizon optimization under real forecasts (see Appendix B as well as C) and approximated MILP solutions: Long-term optimizations are stopped at a $\text{MILP}_{\text{gap}}<0.1\,\%$ or after 30s; short-term optimizations at a $\text{MILP}_{\text{gap}}<0.1\,\%$ or after 10s.
[…]
Figure 9. Additional carbon savings under realistic conditions for Wiki (de) across all regions and $\text{QoR}_{\text{target}}=0.5$.
@@
Outlook of annual emissions.  Long-term optimizations are computationally cheap when accepting near-optimal solutions—within just 30 seconds, the mean $\text{MILP}_{\text{gap}}$ reached 0.53 % accross all experiments. This allows us to generate a broad spectrum of solutions for varying $\text{QoR}_{\text{target}}$ and request forecasts, to represent, for instance, worst-case or best-case scenarios. Figure 10 demonstrates emission outlooks for Wiki (en) traces in California. For simulating best-case and worst-case request forecasts, we use the lower and upper bound of the 80 % prediction intervals provided by Prophet, respectively. Each week, we solve 33 long-term optimizations, covering the three forecasts and $\text{QoR}_{\text{target}}=\{0,0.1,...,1\}$.
(a) Annual emission outlook at three discrete points in time.
(b) Annual emission outlook; weekly for $\text{QoR}_{\text{target}}=0.5$.
Figure 10. Projected emissions (blue) based on long-term forecasts. The interval shows the range between best- and worst-case forecasts. Red indicates emissions already incurred.
@@
Figure 10(b) shows that the initial estimate was overstated, although still in the range of the lower bound It depicts the weekly estimates for annual emissions when aiming for $\text{QoR}_{\text{target}}=0.5$. The upper and lower bound, defined by the worst-case and best-case forecasts, narrow over time as incurred emissions accumulate.
5.5. Automatic QoR Adaptation
Figure 11. Daily QoR for California (CISO), Texas (ERCOT), and Poland (PL). Our approach provides the most consistent QoR throughout the year.
Lastly, we illustrate the automatic adaptation of $\text{QoR}_{\text{target}}$ introduced in Section 4.2 on Wiki (en) in three exemplary regions. For each region, we define an annual budget based on the resulting emissions of the corresponding upper-bound experiment (Section 5.3) for $\gamma=24$ hours and $\text{QoR}_{\text{target}}=0.5$. Consequently, under perfect forecasts, the optimal approach would be to choose $\text{QoR}_{\text{target}}=0.5$ for every time step. We compare our approach to two greedy baselines that have access to the same forecasts as our approach:
@@ 5.5. Automatic QoR Adaptation
    Greedy (constant) divides the annual budget into equally sized hourly budgets for which it solves (15). Any remaining or overspent budget is evenly distributed across the remaining time intervals.
@@
    Greedy (weighted) also solves (15), but weights the hourly budgets based on forecasted load and carbon intensity. Every 24 hours, as request and carbon intensity forecasts are updated, weightings get adjusted.
Figure 11 shows the daily QoR of the greedy baselines and our automatic QoR adaptation. Since Greedy (constant) does not account for fluctuations in requests or carbon intensity, the provided QoR varies significantly throughout the year. Greedy (weighted) delivers more consistent QoR in all experiments but tends to fall back to either $\text{QoR}=1$ (NL) or $\text{QoR}=0$ (CISO and NYISO) toward the end of the year due to forecast errors. In the CISO and NYISO regions, it even exceeds the budget by 0.8 and 1.2 %, respectively. Our approach provides the most consistent service across all experiments as it periodically adjusts $\text{QoR}_{\text{target}}$ based on long-term optimizations. The annual budget is utilized to 98.2–99.8 %.